OpenStudy (anonymous):

-3(y-6)+25=8(7-4y) does y= -13 over 29 ( -13/29)

OpenStudy (anonymous):

I got 33/29

OpenStudy (anonymous):

so what i did was distribute first -3y+18+25=56-32y

OpenStudy (anonymous):

then i combined like terms

OpenStudy (anonymous):

-3y+23=56-32y

OpenStudy (anonymous):

after that, i moved 32y over and 23 over 29y=33

OpenStudy (anonymous):

then i just divided 29 on both sides, effectively getting y=33/29

OpenStudy (anonymous):

what i did : -3(y-6)+25=8(7-4y) -3y+18+25=56-32y -32y+3y= -29y 18+25=43=56-29y -13=29y and then i got -13/29 even though while typing this i realized i messed up .

OpenStudy (anonymous):

But you do see what you did wrong right?

OpenStudy (anonymous):

and you know how to fix it? As long as you realize your mistakes (be they little or big) and are able to fix them, you should do fine

OpenStudy (anonymous):

i still ended up getting a different answer ..

OpenStudy (anonymous):

what answer did you get?

OpenStudy (anonymous):

-20/-29

OpenStudy (anonymous):

actually you might be right, holdup

OpenStudy (anonymous):

so would it be 1.14 ?

OpenStudy (anonymous):

okay .

OpenStudy (anonymous):



OpenStudy (anonymous):

yup , i got that as well :)

OpenStudy (anonymous):

i went wrong combining the like terms

OpenStudy (anonymous):

ooooh , well we all make mistakes here and here ! but i still appreciate you helping me out :) we both learning something and helping eachother out as well! ha ! :)

OpenStudy (anonymous):